please help: solve using any Algebra Method
3x-2y=-5
x-4y=1

Answer by mananth(16946)

3 x -2 y = -5 .............1
1 x -4 y = 1 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-6 x +4 y = 10
1 x -4 y = 1
Add the two equations
-5 x = 11
/ -5
x=11/5
plug value of x in (1)
3 x + -2 y = -5
-6.6 + -2 y = -5
-2 y = -5 + 6.6
-2 y = 1.6
y = - 4/5
-11/5 ,- 4/5
m.ananth@hotmail.ca